OpenStudy (anonymous):

Which values of x will make the absolute value equation true? |x + 1| = 2

OpenStudy (anonymous):

well todays list of questions i got from a website deal with Absolute Value Equations in One Variable

OpenStudy (anonymous):

so yippiii for whom ever likes solving those

OpenStudy (uri):

|x + 1| = 2 x=2-1 x=1 |x + 1| = 2 x+1=-2 x=-2-1 x=-3

OpenStudy (anonymous):

good job uri ;D